OpenStudy (anonymous):

Help Please !! Will Fan and Give Medals !! A system of equations is shown below: 3x+8y=12 2x+2y=3 Part A:create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other.show the steps to do this. Part B:show that the equivalent system has the same solution as the Original system of equations.

OpenStudy (tkhunny):

Have you considered multiplying the second by -4?

OpenStudy (anonymous):

No I havent I really need help on learning how to do this..

OpenStudy (tkhunny):

Okay, multiply the second by -4.

OpenStudy (anonymous):

Multiply 3 by -4?

OpenStudy (tkhunny):

2x+2y=3 Multiply by -4 -8x - 8y = -12

OpenStudy (anonymous):

Does that become the overall equation..Im confused..Sorry..

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 3x+8y=12&\to &\quad 3x\cancel{ +8y }=12\\ 2x+2y=3&{\color{brown}{ \times -4}}\to &-8x\cancel{ -8y }=-12\\ \hline\\ &&\qquad \qquad ? \end{array}\)

OpenStudy (anonymous):

Would it be 3??:))

OpenStudy (jdoe0001):

3? what do you mean?

OpenStudy (anonymous):

The question Mark..Will the Overall answer be 3??:))

OpenStudy (jdoe0001):

hmmm nope, is a sum so you add the values vertically down notice the variables are already lined up vertically

OpenStudy (anonymous):

So the questions already answered thats all i needed to do..Can i ask a question?:)

OpenStudy (jdoe0001):

sure

OpenStudy (anonymous):

Were did you get the -4 from?? :))

OpenStudy (jdoe0001):

is just an arbitrary value used to multiply one of the equations

OpenStudy (jdoe0001):

what did the sum give you?

OpenStudy (anonymous):

It gave me 3 :))

OpenStudy (jdoe0001):

well... is not 3... notice there's a variable.... that you seem to be missing there

OpenStudy (anonymous):

Ohhhh xD is the Missing Variable -4...is that what i was suppossed to find?

OpenStudy (tkhunny):

You are supposed to find values for x and y that make BOTH equations true simultaneously. 1) x+8y=12 2) 2x+2y=3 One method is called elimination. Multiple 2) BY -4 1) x+8y=12 3) -8x-8y=-12 Add 1) and 3), combining like terms. 4) -7x = 0 Solve for x 5) x = 0 Substitute 5) into 1) to find the corresponding value for y. 6) 0+8y=12 Solve for y 7) y = 12/8 = 3/2 Check in 2) 8) 2(0)+2(3/2) = 0 + 3 = 3. Done. The solution is (0,3/2). Perhaps it is more clear to write (x = 0, y = 3/2)

OpenStudy (anonymous):

Thanks so verry Much @tKhunny:)) <3

OpenStudy (tkhunny):

Normally, I would insist that YOU do all the work. However, sometimes a discussion just drags on too long and I want to put a stop to it before someone becomes discouraged. Let's see you do one.